A service station has 36 octane and 60 octane gasoline. How many liters of each gasoline should be mixed to provide 12 liters of 50 octane gasoline for a chemistry experiment?
Answer by stanbon(75887)

A service station has 36 octane and 60 octane gasoline. How many liters of each gasoline should be mixed to provide 12 liters of 50 octane gasoline for a chemistry experiment?
---------
Equation::
octane + octane = octane
--------------------
36x + 60(12-x) = 50*12
36x + 60*12 - 60x = 50*12
-----
24x = 10*12
-----
x = (1/2)10 = 5 liters (amt. of 36 octane needed)
-------------
12-x = 7 liters (amt. of 60 octaine needed)
-------------------
Cheers,
Stan H.
-------------------